Grade 11IIT JEE Entrance Exam

A locomotive of mass m starts moving so that its velocity varies according to law v=a(s)^1/2 where a is a constant and s is the distance covered Find total work performed by all the forces which are acting on the locomotive during the first t seconds after begining the motion?

Jigyasha Sharma
11 Years agoGrade 11

2 Answers

Saurabh Kumar
11 Years ago
Since, it is given , dS/dt =a(s)^1/2, calculate the va;lue of ‘s’ from this relation by integrating with in the limit.
After calciulating ‘s’, put this value in the the given relation to find ‘v’.
similarly, calculate acceleration by differentiating ‘v’.
Put the value in the , W = F . S = m a .s
Put the value and get the answer..........mat4/8....


Good Luck...................................
Adrija Chatterjee
7 Years ago
v=a(s)^1/2
or ds/dt=a(s)^1/2
or ds/(s)^1/2=adt
Integrating both sides we get
2(s)^1/2=at
or s=a^2×t^2/4
W=F×s
=ma×a^2×t^2/4
=ma^3×t^2/4